Ricardo is 4 years less than twice as old as his sister. The sum of their ages is 20. How old is Ricardo?
Answer by ptaylor(2198)

Let x=his sister's age
Then (2x-4)=Ricardo's age
Now we are told that the sum of their ages is 20. So:
x+(2x-4)=20 get rid of parens
x+2x-4=20 add 4 to both sides
x+2x-4+4=20+4 collect like terms
3x=24 divide both sides by 3
x=8--------------------------------his sister's age
2x-4=2*8-4=16-4=12--------------------Ricardo's age

CK
8+12=20
20=20
and
2*8-4=12
12=12


Hope this helps-----ptaylor